OpenStudy (anonymous):

Identify the 25th term of the arithmetic sequence 2, 1 3/5, 1 1/5 … −73/5 −8 −122/5 −13

hartnn (hartnn):

do you know what is an arithmetic sequence ? can you find common difference ?

OpenStudy (anonymous):

yea the common difference can be founded by adding the biggest from the smallest right?

hartnn (hartnn):

not actually, common difference 'd' is the difference between the consecutive terms like d = 2nd term - 1st term = 3rd term - 2nd term = and so on ...

OpenStudy (anonymous):

oh okay okay so id take 2 - 1 3/5 wich equals 2/5

hartnn (hartnn):

actually 2nd term is 1 3/5 and 1st term is 2 so d = 1 3/5 - 2 = -2/5 got this ?

OpenStudy (anonymous):

oh okay yea got it

hartnn (hartnn):

and 1st term = a1 = 2 now plug those in the general formula \(\large a_n = a_1 + (n-1)d\)

OpenStudy (anonymous):

-7 2/3

OpenStudy (anonymous):

so its A.

OpenStudy (anonymous):

THANKS

hartnn (hartnn):

2-48/5 = 10 -48/5 = -38/5 = - 7 3/5 yes, thats correct :) welcome ^_^

OpenStudy (whpalmer4):

It's important to solidify your understanding of the correct meaning of "common difference" (and similarly "common ratio" for geometric sequences) as getting it backwards will lead to many wrong answers...The common difference is the value added to the first term to get the second term, as you can see from the formula \[a_n = a_1 + (n-1)d\]where we find subsequent terms by repeatedly adding the common difference \(d\) to the first term.